OpenStudy (anonymous):

It's me again :):):):) A taxi company charges riders a fixed charge of $2.50 plus $1.50 per mile. How many miles must a rider go to have an average cost per mile of $2.00?

OpenStudy (anonymous):

2.5+1.5x=2x Re-arrange and solve for x 2.5=0.5x, x=5 2.5+7.5=10=2*5. Just to show it solves it.

OpenStudy (anonymous):

you are saving my butt today with math -- thank you! :)... i still don't know if i understand the fraction one i previously posted, though :\